OpenStudy (anonymous):

(14x2 + 7x) + 9x2 = is it 30x5?

OpenStudy (anonymous):

NO ITS 23 x^2+7 x

OpenStudy (anonymous):

\[14x^2+7x+9x^2=23x^2+7x\]

OpenStudy (diyadiya):

noooooooo

OpenStudy (stormfire1):

Since it's all being added you can drop the parentheses and just add like terms: 14x^2 + 9x^2 = 23x^2... 23x^2 + 7x is the answer

OpenStudy (anonymous):

14x^2 + 7x + 9x^2 = 23x^2 + 7x s d answer